# Interpreting Low-level Vision Models with Causal Effect Maps

### v2

@@ II Related Work
Low-level Vision: The LV tasks focus on computer vision tasks where images serve as input, processing, and output units. The primary objective is to reconstruct or enhance low-quality images that suffer from diverse forms of degradation. Representative LV tasks, categorized by the specific type of degradation, encompass super-resolution (SR), denoising (DN), deraining (DR), and more.
@@
In addition to these advancements, it is important to recognize that low-vision (LV) tasks are often intertwined, and low-quality images may suffer from various degradations. To address this complexity, researchers have proposed several general LV models capable of solving multiple tasks within a single network backbone. Notable examples include SWinIR, MPRNet, and XRestormer, which are designed to handle multitasking problems effectively, providing a comprehensive solution for various image restoration and enhancement challenges [1, 2, 4].
@@ IV-A General Intervention Framework
Inspired by the notion of intervention in causality [9], we find that the method of interpreting the network through interventions (or perturbations) dovetails seamlessly with our objective. The intervention serves as a direct and effective method for assessing the existence of a direct causal relationship between variables [26, 9]. When examining the causal relationship between a variable $p_{i}$ from the input set $I$ and its corresponding output $\mathcal{F}(I)$ through the function $\mathcal{F}(\cdot)$, we can disrupt the observed correlation of the original input and output by intervention, denoted as $\textit{do}(p_{i}=x)$. Then we can analyze the subsequent change as follows:
@@
In the context of LV, $p_{i}$ signifies a patch in the input image, $I$ represents the original input image, $\mathcal{F}$ denotes the network, and $\mathcal{M}_{R}$ ( i.e., $\mathcal{M}$ in Eq. 1) signifies the assessment of image quality for a selected ROI $R$. Note that, intervening with patches within the ROI would disrupt their original content, making the analysis meaningless. Thus our focus is on patches outside the ROI. At the same time, since the patches within the ROI are the direct contributors to the ROI restoration, we define the causal effect of those patches as +$\infty$+ ∞. To sum up, th pipeline can be described as Alg. 1.
@@ IV-B Defining LV Interventions
LV encompasses various tasks like super-resolution, denoising, deraining, etc. Degradation represents the inherent characteristic of various tasks and serves as a key differentiator among them. The premise of the LV network’s proper performance is that the input image falls in a degraded distribution that the network has learned before. Therefore, we believe that when LV interventions are performed, the degradation information in the post-intervention image needs to align with the original image. How the model utilizes the image content behind these degradations is our focus.
@@ IV-C Characterizing LV Interventions
However, LV tasks are particularly sensitive to changes in pixel intensity. The abrupt introduction of one simple all-black patch - an anomaly starkly contrasted against its surroundings - can result in a significant departure from the natural image distribution, as depicted in Fig. 4(c). Such an intervention can disrupt the normal behavior of LV networks. This underscores the need to respect the innate characteristics and distribution of an image.
A intuitive method to assess the significance of a variable is to analyze the effect of altering the variable’s contribution, either by increasing or decreasing it from its original state, on the model’s output. For instance, a coffee shop can quantitatively evaluate the impact of the coffee price on sales by observing the changes in sales volume following an increase or decrease in the price, with other factors held constant. In the LV domain, it is challenging to clearly identify the contributions of specific patches and enhance them in a practical way. The ablation of patches, on the other hand, may be easier to achieve and more reliable. Therefore, our intervention strategy is to address the scenario where the variable is in an absence state, presenting an average effect in which the variable has no merit or fault. Building on this understanding, Eq. 1 can be regarded as calculating the causal effect difference between the original state $\mathcal{M}_{R}\mathcal{F}(I)$ and the absence state $\mathcal{M}_{R}(\mathcal{F}(I\mid\text{do}(p_{i}=x)))$.
@@
Figure 5: (a) Intervening all the patches for $T$ times to get a CEM. (b) Dividing the patches into two sets $\mathcal{U}$ and $\mathcal{S}$ in the coarse stage. (c) Refining the causal effect of set $\mathcal{S}$ in the fine stage.
@@ V-B Mechanism Diagnosis
Figure 8: (a) Changing just one pixel in the most negative patch significantly alters the performance of RCAN, whereas (b) the performance change for SwinIR is relatively smaller. These pixels’ RGB values are set to be (0,0,0). They are located in their own most negative patches, respectively.
Fig. 7 compares the LAM and CEM results of SAN, RCAN, and SwinIR, illustrating why causality is important. Analyzing the LAM results, one might assume that these networks are effectively handling the image due to the full image utilization. However, they are actually misled to some extent by the input patches. Specifically, for SAN and RCAN, the most negative patch causes the PSNR to drop by more than 4 dB and 3 dB, respectively. For instance, in the case of RCAN, changing just one pixel in the most negative patch drastically alters the PSNR of the ROI and restores the correct texture direction as shown in Fig. 8. In contrast, the performance change of SwinIR is smaller than that of RCAN. This demonstrates the instability of RCAN when processing this image.
@@ V-C Differences Between LV Tasks
Figure 10: Average percentage of patches with positive and negative causal effects of all the models’ CEMs on the test set is exhibited in the pie chart. The colorbars indicate the average causal effect range of models which are scaled consistently. The proposal years of the models are also recorded.
@@ V-D Interpreting General LV Models
Furthermore, We also conducted experimental analyses on pre-trained Multiple-In-One (MIO) networks trained with a broader range of tasks [54]. Specifically, SRResNet-MIO and SwinIR-MIO were trained with 7 mixed tasks, including super-resolution, deblurring, denoising, deJPEG, deraining, dehazing, and low-light enhancement. The CEM results are presented in Fig. 13, which demonstrate the phenomenon that the MIO models utilize a more localized image area when dealing with SR task.
@@ VI-A Natural Image Library
Figure 14: Different CEMs of SwinIR, RCAN, and SRResNet generated with DIV2K and Flickr2K.
@@ VI-B Patch Size
VI-C Accelrating Strategy
where $x$ and $y$ denote the indices of columns and rows, respectively. $\|\cdot\|_{2}$ is the $\ell_{2}$ norm. It should be noted that the method of calculating the gradient does not need to strictly adhere to the above equation, as long as it serves as an effective measure of variation in the content of the image. Additionally, the hyperparameters for the number of interventions in the coarse and fine stages of CEM calculation can be discussed.

### v3

@@ II Related Work
Low-level Vision: The LV tasks focus on computer vision tasks where images serve as input, processing, and output units. The primary objective is to reconstruct or enhance low-quality images that suffer from diverse forms of degradation [10, 11, 12, 13, 14, 15]. Representative LV tasks, categorized by the specific type of degradation, encompass super-resolution (SR), denoising (DN), deraining (DR), and more.
@@
In addition to these advancements, it is important to recognize that low-vision (LV) tasks are often intertwined, and low-quality images may suffer from various degradations. To address this complexity, researchers have proposed several general LV models capable of solving multiple tasks within a single network backbone. Notable examples include SWinIR [1], MPRNet [2], and X-Restormer [4], which are designed to handle multitasking problems effectively, providing a comprehensive solution for various image restoration and enhancement challenges.
@@ IV-A General Intervention Framework
Inspired by the notion of intervention in causality [9], we find that the method of interpreting the network through interventions (or perturbations) dovetails seamlessly with our objective. The intervention serves as a direct and effective method for assessing the existence of a direct causal relationship between variables [32, 9]. When examining the causal relationship between a variable $p_{i}$ from the input set $I$ and its corresponding output $\mathcal{F}(I)$ through the function $\mathcal{F}(\cdot)$, we can disrupt the observed correlation of the original input and output by intervention, denoted as $\textit{do}(p_{i}=x)$. Then, we can analyze the subsequent change as follows:
@@
In the context of LV, $p_{i}$ signifies a patch in the input image, $I$ represents the original input image, $\mathcal{F}$ denotes the network, and $\mathcal{M}_{R}$ ( i.e., $\mathcal{M}$ in Eq. 1) signifies the assessment of image quality for a selected ROI $R$. Note that, intervening with patches within the ROI would disrupt their original content, making the analysis meaningless. Thus our focus is on patches outside the ROI. At the same time, since the patches within the ROI are the direct contributors to the ROI restoration, we define the causal effect of those patches as +$\infty$+ ∞. To sum up, the pipeline can be described as Alg. 1.
@@ IV-B Defining LV Interventions
LV encompasses various tasks like super-resolution, denoising, deraining, etc. Degradation represents the inherent characteristic of various tasks and serves as a key differentiator among them. The premise of the LV network’s proper performance is that the input image falls in a degraded distribution that the network has learned before. Therefore, we believe that when LV interventions are performed, the degradation information in the post-intervention image needs to align with the original image. Our focus is on how the model utilizes the image content behind these degradations.
@@ IV-C Characterizing LV Interventions
However, LV tasks are particularly sensitive to changes in pixel intensity. The abrupt introduction of one simple all-black patch - an anomaly starkly contrasted against its surroundings - can result in a significant departure from the natural image distribution, as depicted in Fig. 4(c). Such interventions can disrupt the normal behavior of LV networks, leading to significant changes that may render subsequent analyses unreliable.
An intuitive method to assess the significance of a variable is to analyze the effect of altering the variable’s contribution, either by increasing or decreasing it from its original state, on the model’s output. For instance, a coffee shop can quantitatively evaluate the impact of the coffee price on sales by observing the changes in sales volume following an increase or decrease in the price, with other factors held constant. In the LV domain, it is challenging to clearly identify the contributions of specific patches and enhance them in a practical way. The ablation of patches, on the other hand, may be easier to achieve and more reliable. Therefore, our intervention strategy is to address the scenario where the variable is in an absence state, presenting an average effect in which the variable has no merit or fault. Building on this understanding, Eq. 1 can be regarded as calculating the causal effect difference between the original state $\mathcal{M}_{R}\mathcal{F}(I)$ and the absence state $\mathcal{M}_{R}(\mathcal{F}(I\mid\text{do}(p_{i}=x)))$.
@@
Figure 5: (a) Intervening all the patches for $T$ times to get a CEM. (b) Dividing the patches into two sets $\mathcal{U}$ and $\mathcal{S}$ in the coarse stage. (c) Refining the causal effect of set $\mathcal{S}$ in the fine stage. The patches are sampled according to the probability density $\mathcal{D}$, which is further discussed in Sec. VI-C.
@@ V-B Mechanism Diagnosis
Figure 8: (a) Changing just one pixel in the most negative patch significantly alters the performance of RCAN, whereas (b) the performance change for SwinIR is relatively smaller. These pixels’ RGB values are updated to (0, 0, 0) over 2000 iterations. They are located in their own most negative patches, respectively. PSNR Improvement During Pixel Optimization Iterations for (c) RCAN and (d) SwinIR.
Fig. 7 compares the LAM and CEM results of SAN, RCAN, and SwinIR, demonstrating the importance of causality. From the LAM results, one might assume that all these networks effectively handle the image due to their extensive utilization of the input. However, the CEM results further unveil the underlying issues: a significant portion of the input patches have negative effects, misleading the outputs of different networks to varying degrees. For SAN and RCAN, the most negatively contributing patch causes the PSNR to drop by 4.98 dB and 3.43 dB, respectively. In contrast, for SwinIR, the input patches result in at most a 0.31 dB drop. This demonstrates that SAN and RCAN are more vulnerable to being misled by the information in the input image compared to SwinIR.
We next design an experiment to explore whether modifying the most negatively contributing patch can improve the network’s performance and to what extent such improvement can be achieved. In this experiment, a single pixel from the most negative patch is selected and set as a learnable parameter. The loss function is defined as the L2 loss between the output and the ground truth (GT) ROI. Using the Adam optimizer, this pixel is updated over 2000 iterations to enhance the reconstruction performance. In the case of RCAN, modifying just one pixel in the most negative patch significantly increases the PSNR by 5.13 dB and corrects the erroneous texture, as shown in Fig. 8(a). In contrast, for SwinIR, the original input pixel does not severely mislead the reconstruction process. Updating the pixel does not result in a qualitative transformation of the reconstruction, and the performance improvement is smaller compared to RCAN, as illustrated in Fig. 8(b). Moreover, as shown in Fig. 8(c) and (d), RCAN’s performance is more sensitive to pixel modifications and is more significantly improved by such updates compared to SwinIR. This further demonstrates that RCAN is more susceptible to the influence of input information, with the original pixel contributing to greater negative effects on its reconstruction process.
@@ V-C Differences Between LV Tasks
Figure 10: Average percentage of patches with positive and negative causal effects of all the models’ CEMs on the test set is exhibited in the pie chart. The colorbars indicate the average causal effect range of models, which are scaled consistently. The proposal years of the models are also recorded.
@@ V-D Interpreting General LV Models
Furthermore, We also conducted experimental analyses on pre-trained Multiple-In-One (MIO) networks trained with a broader range of tasks [57]. Specifically, SRResNet-MIO and SwinIR-MIO were trained with 7 equally mixed tasks, including super-resolution, deblurring, denoising, deJPEG, deraining, dehazing, and low-light enhancement. The CEM results are presented in Fig. 13, which demonstrate the phenomenon that the MIO models utilize a more localized image area when dealing with SR task.
@@ VI-A Natural Image Library
Figure 14: Different CEMs of SwinIR, RCAN, and SRResNet generated with patches sampled from DIV2K and Flickr2K.
@@ VI-B Patch Size
VI-C Accelrating Strategy
where $x$ and $y$ denote the indices of columns and rows, respectively. $\|\cdot\|_{2}$ is the $\ell_{2}$ norm.
Figure 17: Several visual examples of patches with $G$ values sampled from the following ranges are presented: (a) 0–128, (b) 128–256, (c) 256–384, and (d) 384–512. The corresponding $G$ values are indicated above each patch.
Several samples from different ranges of $G$ are presented in Fig. 17. We observe that patches with smaller $G$ values tend to have simpler and smoother content, while patches with larger $G$ values exhibit more complex structures and greater pixel-to-pixel differences. Additionally, in the probability density $\mathcal{{D}}$, patches sampled with smaller $G$ values are more frequent, which aligns with intuition—flat regions are more prevalent than regions with significant variations in natural images. It should be noted that the method of calculating the gradient does not need to strictly adhere to the above equation, as long as it serves as an effective measure of variation in the content of the image. Additionally, the hyperparameters for the number of interventions in the coarse and fine stages of CEM calculation can be discussed.